Project UMUC is to produce 200 widgets and is scheduled to take five weeks. Each unit...

  1. Project UMUC is to produce 200 widgets and is scheduled to take five weeks. Each unit is planned to cost $90. The project is severely cost constrained. Performance data for the project at the end of week three is presented below:
  • 120 total units were planned to be produced
  • 130 units have actually been produced
  • The financial manager reported that the business had actually spent $13,000 on the project by the end of week three.

Answer the following questions; show all work:

  1. What is the forecast of project cost at completion assuming current cost performance efficiency remains the same? How much budget variance is expected at completion?
  2. What is the forecast of funding needed to complete the project (from this point forward)?
  3. What cost performance efficiency would be required for the remainder of the project to complete the project within the original budget?
  4. As the project financial manager, what recommendations would you make?
Answer #1

Please rate the earlier solution where first four parts were answered.

What is the forecast of project cost at completion assuming current cost performance efficiency remains the same? How much budget variance is expected at completion?

Forecast of project cost at completion = Actual cost / Actual number of units x total number of units = (13,000 / 130) x 200 = $ 20,000

Budget variance expected at completion = Forecast of project cost at completion - budget cost = 20,000 - 200 x 90 = $ 2,000

What is the forecast of funding needed to complete the project (from this point forward)?

forecast of funding needed to complete the project (from this point forward) = Forecast of project cost at completion - cost incurred so far = 20,000 - 13,000 = 7,000

What cost performance efficiency would be required for the remainder of the project to complete the project within the original budget?

Balance budget cost = (Budget cost - cost incurred so far) =(200 x 90 - 13,000) = 5,000

Balance number of units = (200 - 130) = 70

Cost performance efficiency = Earned Value on balance units / Balance budget cost = 70 x 90 / 5,000 = 126.0%

As the project financial manager, what recommendations would you make?

Recommendations:

  • Cost control: The budget has exceed so far and is likely to remain exceeded unless cost is controlled. hence, focus on cost control.
  • Arrange for external funding, the budget is likely to be exceeded. In order to prevent time overrun, look for funding right now.